# EntropyStop: Unsupervised Deep Outlier Detection with Loss Entropy

### v2

@@ Abstract.
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining; August 2024; Barcelona, Spainisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Anomaly detectionccs: Computing methodologies Neural networksccs: Computing methodologies Machine learning
@@ 2.2. Early Stopping Techniques
Early stopping is an effective and broadly used technique in machine learning. Early stopping algorithms are designed to monitor and stop the training when it no longer benefits the final performance. A well-known application of early stopping is to use it as a regularization method to tackle overfitting problems with cross-validation, which can be traced back to the 1990s (Morgan and Bourlard, 1989). Recently, with a deeper understanding of learning dynamics, early stopping is also found practical in noisy-labeled scenarios (Li et al., 2020; Bai et al., 2021; Xia et al., 2020; Arpit et al., 2017). According to these previous studies, overfitting to the noisy samples in the later stage of training decreases the model’s performance, and can be mitigated by early stopping. Previous works show the outstanding ability of early stopping to deal with noisy learning environments. However, existing researches focus on supervised or semi-supervised settings, while early stopping in unsupervised contaminated training set is significantly more challenging.. To our best knowledge, we are the first to apply a label-free and distribution-based heuristic to explore the potential of early stopping in Unsupervised OD on contaminated training sets.
@@ 5.2.2. Efficiency Result
(11) $Time_{train}(M)=\frac{1}{|\mathcal{D}|}\sum_{D\sim\mathcal{D}}\frac{\text{training time}(M,D)}{\text{training time}(VanillaAE,D)}$
where $D$ represents one of the 47 datasets, $\mathcal{D}$ denotes the collection of all 47 datasets, and $M$ signifies any model among VanillaAE, EntropyAE, RandNet, and ROBOD. For the training time calculation, we ensure that all models have the same batch size of 64 and number of epochs of 250 to guarantee identical iteration counts. $Time_{train}(M)$ reveals the relative training time required compared to VanillaAE. Through this metric, we observe that, compared to VanillaAE, ROBOD, and RandNet, EntropyAE only requires under 8%, 2%, and 0.3% of the training time, respectively. This demonstrates the effectiveness of early stopping in saving time. The detailed comparison result on efficiency are list in Table 2.
@@
Train Time 1 0.077 23.05 3.51
@@ 5.3. Negative Correlation Study
All the AUC and $H_{L}$ curves are shown in Fig. 8, 9, 10, 11 (AE) and Fig. 12, 13, 14, 15 (DeepSVDD) in Appx. 6 to demonstrate the negative correlation between the two. The distribution of Pearson correlation coefficient values across 47 datasets are shown in Fig. 6. These results show that while $H_{L}$ has a strong negative correlation with AUC on more than half of the 47 datasets, the remaining datasets show a weak or even positive correlation. Basically, the reason for invalidity can be attributed to the following aspects:
@@ A.2.1. Assumptions
$P(\delta+\Delta\delta>0)>P(\delta>0)\rightarrow P(\Delta\delta>0)>0.5$
$P(\delta+\Delta\delta>0)>P(\delta>0)\rightarrow P(\Delta\delta>0)<0.5$
Here, we assume that if the $AUC$ increases (or decreases) after a single gradient update, the decrease in outliers’ losses is more (or less) likely to be smaller than the decrease in inliers’ losses.
@@ Appendix C Limitation analysis and case study
In the experiment described in Sec 6, we assess the negative correlation between loss entropy $H_{L}$ and AUC utilizing the Pearson correlation coefficient, abbreviated as $r$ for clarity. Fig. 8, 9, 10, and 11 illustrate the evolution of AUC and entropy curves of AE throughout the training period for 47 datasets, ranked by descending order of their negative correlation strength. Notably, while the loss entropy $H_{L}$ demonstrates a strong negative correlation with AUC across several datasets, there are still some datasets that exhibit weak or even positive correlations, such as ALOI. We attribute this primarily to the following two reasons. We categorized the datasets in Table 6 on which entropy stop does not perform well.
@@ D.2.3. Conclusion
Table 11. Training Time (Compared to VanillaAE) of four AE models on 47 datasets

### v3

@@ Abstract.
copyright: acmlicensedjournalyear: 2024conference: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining; August 25–29, 2024; Barcelona, Spainbooktitle: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining (KDD ’24), August 25–29, 2024, Barcelona, Spaindoi: 10.1145/3637528. 3671943isbn: 979-8-4007-0490-1/24/ 08ccs: Computing methodologies Anomaly detectionccs: Computing methodologies Neural networksccs: Computing methodologies Machine learning
@@ 2.2. Early Stopping Techniques
Early stopping is an effective and broadly used technique in machine learning. Early stopping algorithms are designed to monitor and stop the training when it no longer benefits the final performance. A well-known application of early stopping is to use it as a regularization method to tackle overfitting problems with cross-validation, which can be traced back to the 1990s (Morgan and Bourlard, 1989). Recently, with a deeper understanding of learning dynamics, early stopping is also found practical in noisy-labeled scenarios (Li et al., 2020; Bai et al., 2021; Xia et al., 2020; Arpit et al., 2017). According to these previous studies, overfitting to the noisy samples in the later stage of training decreases the model’s performance, and can be mitigated by early stopping. Previous works show the outstanding ability of early stopping to deal with noisy learning environments. However, existing researches focus on supervised or semi-supervised settings, while early stopping in unsupervised contaminated training set is significantly more challenging. To our best knowledge, we are the first to apply a label-free and distribution-based heuristic to explore the potential of early stopping in Unsupervised OD on contaminated training sets.
@@ 5.2.2. Efficiency Result
(11) Average Train $\text{{Average Train Time}}(M)=\frac{1}{|\mathcal{D}|}\sum_{D\sim\mathcal{D}}\frac{\text{training time}(M,D)}{\text{training time}(VanillaAE,D)}$
(12) Total Train $\text{{Total Train Time}}(M)=\frac{\sum_{D\sim\mathcal{D}}\text{training time}(M,D)}{\sum_{D\sim\mathcal{D}}\text{training time}(VanillaAE,D)}$
where $D$ represents one of the 47 datasets, $\mathcal{D}$ denotes the collection of all 47 datasets, and $M$ signifies any model among VanillaAE, EntropyAE, RandNet, and ROBOD. We ensure that all models have the same batch size of 64 and number of epochs of 250 to guarantee identical iteration counts. Train $\text{{Train Time}}(M)$ reveals the average relative training time required compared to VanillaAE while Total Train $\text{{Total Train Time}}(M)$ reveals the total time required compared to VanillaAE. In Table 2, we observe that, compared to VanillaAE, ROBOD, and RandNet, EntropyAE only requires under 8%, 2%, and 0.3% of the average training time, respectively. For the total training time, the advantage of $EntropyAE$ are more significant. This demonstrates the effectiveness of early stopping in saving time. The detailed comparison result on efficiency are list in Table 2.
@@
Average Train Time 1 0.077 23.05 3.51
Total Train Time 1 0.01 35.03 4.02
Figure 6 displays the time required by EntropyAE across 47 datasets. The early stopping mechanism is more effective on larger datasets, as they contain more batches per epoch, resulting in more iterations. In some large datasets, EntropyAE stops training before completing a single epoch.
Figure 6. The relative time (compared to VanillaAE) taken by EntropyAE across different dataset sizes.
@@ 5.3. Negative Correlation Study
All the AUC and $H_{L}$ curves are shown in Fig. 9, 10, 11, 12 (AE) and Fig. 13, 14, 15, 16 (DeepSVDD) in Appx. 6 to demonstrate the negative correlation between the two. The distribution of Pearson correlation coefficient values across 47 datasets are shown in Fig. 7. These results show that while $H_{L}$ has a strong negative correlation with AUC on more than half of the 47 datasets, the remaining datasets show a weak or even positive correlation. Basically, the reason for invalidity can be attributed to the following aspects:
@@ 6. Conclusion
Acknowledgement
This work is supported by the National Science and Technology Major Project 2021ZD0114501.
@@ A.2.1. Assumptions
$E(\delta)>0,P(\delta+\Delta\delta>0)>P(\delta>0)\rightarrow P(\Delta\delta>0)>0.5$
$E(\delta)>0,P(\delta+\Delta\delta>0)<P(\delta>0)\rightarrow P(\Delta\delta>0)<0.5$
Here, we assume that if the loss gap exists and the $AUC$ increases (or decreases) after a single gradient update, the decrease in outliers’ losses is more (or less) likely to be smaller than the decrease in inliers’ losses.
@@ Appendix C Limitation analysis and case study
In the experiment described in Sec 7, we assess the negative correlation between loss entropy $H_{L}$ and AUC utilizing the Pearson correlation coefficient, abbreviated as $r$ for clarity. Fig. 9, 10, 11, and 12 illustrate the evolution of AUC and entropy curves of AE throughout the training period for 47 datasets, ranked by descending order of their negative correlation strength. Notably, while the loss entropy $H_{L}$ demonstrates a strong negative correlation with AUC across several datasets, there are still some datasets that exhibit weak or even positive correlations, such as ALOI. We attribute this primarily to the following two reasons. We categorized the datasets in Table 6 on which entropy stop does not perform well.
@@ D.2.3. Conclusion
Table 11. Average Training Time (Compared to VanillaAE) of four AE models on 47 datasets